OpenStudy (anonymous):

Find the slope of line passing through the points (11, -2) and (2, -2). Would the answer be 9?

Nnesha (nnesha):

nope thats not right would you like to show your work so i can figure it out where you did mistake

OpenStudy (anonymous):

Yeah. That's what I got, but I'm not sure if I'm even doing it right.

Nnesha (nnesha):

okay that;s fine i guess you was doing it right but at the end you did some mistakes its much better if you show your work :) so i can see where you did mistake :)

OpenStudy (anonymous):

Someone told me to figure it out by Y-y1 over x-x1 So -2 - -2 11 - 2 I got 0/9

OpenStudy (danjs):

haha, was that me?

OpenStudy (danjs):

you are right, the y values are the same, it will be horizontal line of slope 0

Nnesha (nnesha):

okay so what is 0/9 equal to

Nnesha (nnesha):

0/9 = 0 so slope is 0 but if you have 9/0 then slope should be undefined you cannot divide by 0

OpenStudy (danjs):

remember the two special cases i typed out before #/0 = undefined slope ,vertical line 0/# - zero slope, horizontal line

OpenStudy (danjs):

lol we are thinking in sync Nnesha.. sorry to interrupt here

OpenStudy (anonymous):

Yeah. So for example if I got 3/8 the answer would be 3?

OpenStudy (danjs):

no, the answer would be slope = 3/8

OpenStudy (anonymous):

And so if it's 0/9 it's undefined because it's 0?

OpenStudy (danjs):

everything else is a valid slope, any fraction the two special cases are vertical and horizontal lines vertical lines being undefined slope horizontal lines being zero slope

OpenStudy (danjs):

no, zero over a number is zero

OpenStudy (anonymous):

Okay I understand it now.

Nnesha (nnesha):

if you divide something by 0 then undefined

OpenStudy (danjs):

\[\frac{ 0 }{ number } = 0~~~horizontal ~slope\] \[\frac{ number }{ 0 }= undefined ~slope\]

OpenStudy (danjs):

anything else, that doesnt include a zero, is a valid slope for the line